# Proto-LeakNet: Towards Signal-Leak Aware Attribution in Synthetic Human Face Imagery

### v1

@@ Abstract
The growing sophistication of synthetic image and deepfake generation models has turned source attribution and authenticity verification into a critical challenge for modern computer vision systems. Recent studies suggest that diffusion pipelines unintentionally imprint persistent statistical traces, known as signal leaks, within their outputs, particularly in latent representations. Building on this observation, we propose Proto-LeakNet, a signal-leak-aware and interpretable attribution framework that integrates closed-set classification with a density-based open-set evaluation on the learned embeddings, enabling analysis of unseen generators without retraining. Operating in the latent domain of diffusion models, our method re-simulates partial forward diffusion to expose residual generator-specific cues. A temporal attention encoder aggregates multi-step latent features, while a feature-weighted prototype head structures the embedding space and enables transparent attribution. Trained solely on closed data and achieving a Macro AUC of 98.13%, Proto-LeakNet learns a latent geometry that remains robust under post-processing, surpassing state-of-the-art methods, and achieves strong separability between known and unseen generators. These results demonstrate that modeling signal-leak bias in latent space enables reliable and interpretable AI-image and deepfake forensics. The code for the whole work will be available upon submission.
@@
Figure 1: Proto-LeakNet: given an input image $x$, latent features are extracted from the pretrained Stable Diffusion 2.1 VAE in Block 1 (Latent Feature Construction). For each diffusion step $t\!\in\!\{0,5,10\}$, the latent $z_{t}$ is concatenated with its low-frequency DCT map $\mathrm{DCT}_{\text{low}}(z_{t})$ to form $f_{t}$, preserving generator-specific structural cues. In Block 2 (Discriminative Backbone), each $f_{t}$ is encoded by a ResNet-18, producing embeddings $\{h_{t}\}$ that are temporally aggregated through the Attention Temporal Pooling module to yield a single representation $\bar{h}$. Block 3 (Prototype-Based Attribution) computes distances between $\bar{h}$ and class prototypes $p_{c,m}$, modulated by a feature-wise gating vector $w$ obtained from a small MLP. The resulting attention-weighted distances are aggregated via a LogSumExp scoring function to produce class probabilities $\pi_{c}$. Symbols: “+” denotes the weighted sum over attention coefficients across timesteps, and “*” indicates the element-wise product between the encoder output and the gating vector.
@@ 1 Introduction
The rapid progress of generative models has transformed digital content creation, enabling the synthesis of highly realistic images and videos, also called deepfakes, that are often indistinguishable from authentic ones [goodfellow2014generative, ho2020denoising, rombach2022high]. New biases such as ‘impostor bias’ [CASU2024301795] further complicate deepfake detection. While these advances have fostered creativity and accessibility, they have also blurred the boundary between real and artificial content, posing serious challenges to media forensics and public trust. As deepfakes proliferate across social, political, and creative domains, distinguishing and attributing their origin has become critical for security, accountability, and digital evidence validation [amerini2025deepfake, verdoliva2020media]. Early research in multimedia forensics focused primarily on classifying whether an image is real or generated [zhang2019detecting, wang2021fakespotter]. However, a deeper forensic question lies in attribution, namely identifying which generative model produced a given image. This task, essential for tracing provenance and assessing responsibility, remains extremely challenging, especially in open-set conditions where unknown generators appear at test time [khoo2022deepfake, bindini2024tiny]. Recent studies have shown that even advanced detectors struggle to generalize beyond the closed domain or to maintain interpretability when facing unseen architectures [corvi2023detection, wang2025openset]. Diffusion models have recently equaled most generative models in terms of image quality, yet they introduce subtle statistical artifacts in their latent representations, known as signal leaks, caused by residual low-frequency information that survives the denoising process [Everaert2024SignalLeak]. These traces, although imperceptible, encode model-specific biases and can serve as reliable forensic cues for source attribution. While prior works have improved detection accuracy, they typically lack robustness to domain shifts and offer limited interpretability, especially under open-set or heavily post-processed conditions. To overcome these issues, we aim to exploit the intrinsic statistical biases embedded in diffusion latents as stable, model-specific signatures. In detail, Proto-LeakNet encodes image $x$ through Stable Diffusion latents and temporal attention, aggregates timestep embeddings via ResNet-18, and performs attribution using prototype-based distances modulated by per-feature attention and gating. This full pipeline can be observed in Figure 1. Our main contributions are the following:
  • We introduce Proto-LeakNet, a signal-leak-aware and interpretable attribution framework that operates entirely in the latent domain of diffusion models, learning generator-specific biases as stable forensic cues.
@@
  • We develop a density-based open-set evaluation using kernel density estimation on the learned embeddings to assess separability between known and unseen generators without retraining.
  • We demonstrate that modeling signal-leak bias in latent space leads to robust attribution under heavy post-processing and provides transparent prototype-level explanations.
This work is organized as follows: Section 2 presents the current state-of-the-art and the works we will be comparing ourselves against. In Section 3 we explain in detail our proposed methodology. Section 4 focuses on explaining the employed dataset and all its specifications. Section 5 presents the results of our experiments and the various ablations we performed. In Section 6 we discuss the results obtained and the limitation of our framework. Finally, Section 7 we evaluate our work and propose future upgrades to enhance the proposed pipeline, establishing where it places itself among the literature.
@@ 2 Related Work
Research on synthetic image detection and attribution has rapidly evolved with the growing realism of generative models. Early detectors focused on binary real-vs-fake discrimination through spatial or frequency inconsistencies. Some early works have established strong baselines that served as foundations for the topic of deepfake detection and attribution. For example Rössler et al. [rössler2019faceforensics] introduced FaceForensics++, an in-depth benchmark for deepfake detection with various image forgery tools. While Nguyen et al. [nguyen2018capsuleforensics] proposes Capsule-Forensics, a method that uses a capsule network to detect various kinds of image modifications and forgeries. Recent works have brought deepfake detection and attribution to the video domain, like StyleFlow (Choi et al. [styleflow2024]) which leverages temporal dynamics of style latent vectors for video deepfake detection. Works like FreqNet (Tan et al. [freqnet2023]) introduced frequency-space learning within CNNs to extract source-agnostic cues from amplitude and phase spectra, while SuSy (Bernabeu-Pérez et al. [susy2024]) provided a large-scale benchmark highlighting the fragility and dataset bias of current detectors under real-world deployment. Recent efforts have shifted toward explicit source attribution. LatentTracer (Wang et al. [latenttracer2024]) performs inversion-based tracing of latent generative models without artificial watermarks, whereas OCC-CLIP (Liu et al. [occlip2024]) reformulates attribution as few-shot one-class classification using CLIP embeddings, enabling model recognition even with limited samples. Alternative approaches aim to generalize by modeling intrinsic generation artifacts: NPR (Tan et al. [npr2024]) captures local pixel correlations induced by up-sampling operations. More recent works have started to employ diffusion latents for synthetic images detection like LATTE [vasilcoiu2025latte] which introduces a transformer-based architecture operating directly in the latent space of diffusion models to detect generated content. By modeling long-range dependencies among latent tokens, it effectively captures generator-specific artifacts. Finally, Everaert et al. [Everaert2024SignalLeak] identified a signal-leak bias in diffusion latents low-frequency mismatches that encode generator-specific traces. Building on this insight, our work introduces Proto-LeakNet, a framework that explicitly models and exploits these latent residuals for robust and interpretable generator attribution under both closed- and open-set conditions. Unlike LATTE, which focuses on discriminative detection through self-attention among latent tokens, Proto-LeakNet explicitly structures the latent space through prototype supervision and temporal attention across multiple diffusion steps. This design aggregates generator-specific cues over time, enabling interpretable attribution rather than binary detection. While LATTE captures static latent correlations, our approach leverages the temporal evolution of signal-leak residuals to model generator-specific biases with higher forensic transparency and open-set separability.
@@ 3 Proposed Method
This section presents the architecture of the proposed Proto-LeakNet framework, divided into three functional blocks. The first, Latent Feature Construction, extracts informative representations from the diffusion latents of Stable Diffusion 2.1. The second, Discriminative Backbone, aggregates these multi-scale features through a temporal attention encoder. Finally, the Prototype-Based Attribution Head interprets the learned representation using class-specific prototypes for interpretable generator attribution.
@@ 3.1 Latent Feature Construction
We extract features directly from the latent domain of Stable Diffusion 2.1 (SD2.1) [rombach2022high]. Each image $x\in\mathbb{R}^{3\times H\times W}$ is resized to $256\times 256$ and encoded through the pretrained SD2.1’s Variational Autoencoder (VAE) into a latent $z_{0}$ scaled by the constant $s=0.18215$, as in
@@
where $\mathcal{E}(\cdot)$ is the VAE encoder. To reveal residual diffusion traces, we reapply the forward diffusion process at discrete steps $t\in\mathcal{T}=\{0,5,10\}$, sampling
@@
where $K$ is the number of diffusion steps and $(\alpha_{t},\sigma_{t})$ follow the cosine schedule with
@@
Each $z_{t}$ is normalized by $\sigma_{t}$ to maintain consistent scale, producing $\{z_{t}/\sigma_{t}\}_{t\in\mathcal{T}}$. To preserve low-frequency structural information, each latent $z_{t}$ is augmented with a low-frequency Discrete Cosine Transform (DCT) map $\mathrm{DCT}_{\text{low}}(z_{t})$ containing the first $8\times 8$ coefficients per channel. The resulting feature tensor is defined as
$f_{t}=\mathrm{concat}\!\left[z_{t},\,\mathrm{DCT}_{\text{low}}(z_{t})\right],$ (4)
where each $f_{t}$ is standardized per channel using statistics computed from the training set. These tensors $\{f_{t}\}$ constitute the multi-scale latent features provided to the temporal attention module.
@@ 3.2 Discriminative Backbone
The feature tensors $\{f_{t}\}_{t\in\mathcal{T}}$ are encoded by a ResNet18 [he2015deepresidual] backbone $\phi(\cdot;\theta)$, with the first convolution adapted to match the channel dimensionality of $f_{t}$. For each timestep $t$, the backbone produces a latent embedding
$h_{t}=\phi(f_{t};\,\theta)\in\mathbb{R}^{D},$ (5)
where $\theta$ denotes the trainable backbone parameters and $D$ is the embedding dimensionality. A learnable attention module assigns relevance to each timestep via
@@ 3.3 Prototype-Based Attribution Head
Each class $c\!\in\!\{1,\dots,C\}$ is represented by $M$ learnable prototypes $p_{c,m}\!\in\!\mathbb{R}^{D}$, which serve as representative points in latent space. Empirically, four prototypes yielded compact yet well-separated latent clusters. An attention feature-wise gating vector $w\!\in\!(0,1)^{D}$ is computed using a small MLP:
@@
where $\tau>0$ determines the aggregation smoothness. Posterior probabilities are computed as
@@
where $B$ is the mini-batch size and $y^{(b)}$ are ground-truth labels. All parameters $\{\theta,A,b,p_{c,m}\}$ are optimized jointly with Adam and weight decay. The main training pipeline can be visualized in full detail in Figure 1.
@@
The most activated prototype $\arg\max_{m}\pi_{c,m}$ identifies the latent region that best matches the input. Together, the feature gates $w$, temporal weights $\{a_{t}\}$, and prototype responsibilities $\{\pi_{c,m}\}$ provide a three-level interpretability hierarchy visible in Figure 2 , describing which prototype, which features, and which diffusion steps drive each attribution decision from layer, as shown in Figure 3.
Figure 2: Plot of the final clusters shaped by the prototypes and attention mechanism.
Figure 3: Plot showing the evolution from layer to layer of the clusters, displaying how separation is emphasized due to prototypes.
@@ 3.4 Representation-Level Generalization via Density Estimation
$p_{\text{KDE}}(h)=\frac{1}{N_{\mathrm{c}}(2\pi\sigma^{2})^{\frac{D}{2}}}\sum_{i=1}^{N_{\mathrm{c}}}\exp\!\left(-\frac{\|h-\bar{h}^{(\mathrm{c})}_{i}\|^{2}}{2\sigma^{2}}\right),$ (16)

### v2

@@ Abstract
The growing sophistication of synthetic image and deepfake generation models has turned source attribution and authenticity verification into a critical challenge for modern computer vision systems. Recent studies suggest that diffusion pipelines unintentionally imprint persistent statistical traces, known as signal-leaks, within their outputs, particularly in latent representations. Building on this observation, we propose Proto-LeakNet, a signal-leak-aware and interpretable attribution framework that integrates closed-set classification with a density-based open-set evaluation on the learned embeddings, enabling analysis of unseen generators without retraining. Acting in the latent domain of diffusion models, our method re-simulates partial forward diffusion to expose residual generator-specific cues. A temporal attention encoder aggregates multi-step latent features, while a feature-weighted prototype head structures the embedding space and enables transparent attribution. Trained solely on closed data and achieving a Macro AUC of 98.13%, Proto-LeakNet learns a latent geometry that remains robust under post-processing, surpassing state-of-the-art methods, and achieves strong separability both between real images and known generators, and between known and unseen ones. The codebase will be available after acceptance.
@@
Figure 1: Proto-LeakNet: given an input image $x$, latent features are extracted from the pretrained Stable Diffusion 2.1 Variational Autoencoder (VAE) in Block 1 (Latent Feature Construction). For each diffusion step $t\!\in\!\{0,5,10\}$, we extract a $z_{t}$ latent which gets normalized with dimension $(4,32,32)$. In Block 2 (Discriminative Backbone), each $z_{t}$ is encoded by a ResNet-18, producing embeddings $\{h_{t}\}$ that are temporally aggregated through the Attention Temporal Pooling module to yield a single representation $\bar{h}$. Block 3 (Prototype-Based Attribution) computes distances between $\bar{h}$ and class prototypes $p_{c,m}$, modulated by a feature-wise gating vector $w$ obtained from a small MLP. The resulting attention-weighted distances are aggregated via a LogSumExp scoring function to produce class probabilities $\pi_{c}$. Symbols: “+” denotes the weighted sum over attention coefficients across timesteps, and “*” indicates the element-wise product between the encoder output and the gating vector.
@@ 1 Introduction
The rapid progress of generative models has transformed digital content creation, enabling the synthesis of highly realistic images and videos, also called deepfakes, that are often indistinguishable from authentic ones [goodfellow2014generative, ho2020denoising, rombach2022high]. While these advances have fostered creativity and accessibility, they have also blurred the boundary between real and artificial content, posing serious challenges to media forensics and public trust. As deepfakes proliferate across social, political, and creative domains, distinguishing and attributing their origin has become critical for security, accountability, and digital evidence validation [xie2025deepfake, verdoliva2020media]. Early research in multimedia forensics focused primarily on classifying whether an image is real or generated [guarnera2020deepfake, wang2021fakespotter]. However, a deeper forensic question lies in attribution, namely identifying which generative model produced a given image. Such task, essential for tracing provenance and assessing responsibility, remains extremely challenging, especially in open-set conditions where unknown generators appear at test time [khoo2022deepfake, bindini2024tiny]. Recent studies have shown that even advanced detectors struggle to generalize beyond the closed domain or to maintain interpretability when facing unseen architectures [corvi2023detection, wang2025openset]. Diffusion models have recently equaled most generative models in terms of image quality, yet they introduce subtle statistical artifacts in their latent representations, known as signal-leaks, caused by residual low-frequency information that survives the noising process [Everaert2024SignalLeak]. These traces, although imperceptible, encode model-specific biases and can serve as reliable forensic cues for source attribution. While prior works have improved detection accuracy, they typically lack robustness to domain shifts and offer limited interpretability, especially under open-set or heavily post-processed conditions. To overcome these issues, we aim to exploit the intrinsic statistical biases embedded in diffusion latents as stable, model-specific signatures. In detail, we propose Proto-LeakNet which encodes image $x$ through Stable Diffusion latents and temporal attention, aggregates timestep embeddings via ResNet-18, and performs attribution using prototype-based distances modulated by per-feature attention and gating. The pipeline is synthetically sketched in Fig. 1. Our main contributions are the following:
  • We introduce Proto-LeakNet, a signal-leak-aware and interpretable attribution framework that operates entirely in the latent domain of diffusion models, learning generator-specific biases as stable forensic cues which demonstrates that modeling signal-leak bias in latent space leads to robust attribution
@@
  • We develop a density-based open-set evaluation using kernel density estimation on the learned embeddings to assess separability between real images and known generators, and between known and unseen ones without retraining.
This work is organized as follows. Section 2 reviews the relevant state of the art. Section 3 introduces our proposed framework. Section 4 describes the dataset and metrics used in our study. Section 5 reports the experimental results and ablations. Section 6 examines the implications and limitations of our approach. Finally, Section 7 summarizes the key findings and outlines future research directions.
@@ 2 Related Work
Beyond binary detection of real vs. fake content [fu2025exploring, cai2025adaptive], recent research has shifted toward identifying the source of generated imagery. Khoo et al. [khoo2022deepfake] outlined the foundations of attribution and categorized deepfake types, identity swap, attribute and expression manipulation, and full-face synthesis, emphasizing interpretability over detection accuracy. Yu et al. [yu2021artificial] proposed a proactive strategy that embeds artificial fingerprints into training data, allowing reliable tracing of model provenance independent of visual artifacts. However, such closed-set settings struggle with unseen generators. Sun et al. [sun2023contrastive] addressed that with the Open-World DeepFake Attribution (OW-DFA) benchmark and the Contrastive Pseudo Learning framework, combining global–local feature voting and confidence-based pseudo-labeling for open-world attribution. Complementary work by Baxevanakis et al. [baxevanakis2025deepfake] showed that standard attribution models generalize poorly across datasets, while contrastive objectives yield more transferable manipulation-aware embeddings. Recent works, for instance, by Tan et al. [freqnet2023] introduce FreqNet, a frequency-domain deepfake detector that forces the model to learn high-frequency features and even operates part of its network in the FFT domain. It includes a dedicated frequency branch between FFT and iFFT to learn source-agnostic cues. Bernabeu-Pérez et al. [susy2024] proposed SuSy, a A CNN-based patch-level forensic classifier that detects whether an image is AI-generated and even attributes it to one of several generator class. It uses a spatial focus: extracting high-contrast patches (via GLCM texture analysis) to train on local artifacts. Wang et al. [latenttracer2024] introduces LatentTracer, a reverse-engineering approach that inverts images into a generative model’s latent space to see if they “belong” to that model. By comparing reconstruction losses, it identifies the source generator without any embedded watermark. Liu et al. [occlip2024] proposes OCC-CLIP, a few-shot one-class attribution framework built upon CLIP embeddings to recognize unseen generators with limited samples. Tan et al. [npr2024] introduces NPR, a method that explicitly targets the telltale upsampling artifacts left by CNN-based generators. It computes local pixel-wise differences (residuals) in small image patches to capture the subtle correlations introduced by upsampling layers. Vasilcoiu et al. [vasilcoiu2025latte] presents LATTE, a transformer-based framework operating directly in the latent space of diffusion models, detecting generated content by modeling long-range dependencies among latent tokens. Finally, Everaert et al. [Everaert2024SignalLeak] identify a signal-leak bias in diffusion latents, showing that low-frequency mismatches encode generator-specific traces. Existing methods still face key limitations, including static latent representations and limited interpretability, often failing to generalize when exposed to unseen generators. Building on these insights, we introduce Proto-LeakNet, a framework that explicitly models and aggregates latent residuals across diffusion steps through prototype supervision and temporal attention. Our approach overcomes these issues by capturing the temporal evolution of generator-specific cues, enabling robust and interpretable attribution under both closed- and open-set conditions.
@@ 3 Proposed Method
The proposed Proto-LeakNet framework is deployed into three functional blocks. Latent Feature Construction, extracts informative representations from the diffusion latents of Stable Diffusion 2.1 [rombach2022high]. Discriminative Backbone, aggregates these multi-scale features through a temporal attention encoder. The Prototype-Based Attribution Head interprets the learned representation using class-specific prototypes for interpretable generator attribution. Finally, once defined the pipeline we introduce the generalization methodology.
@@ 3.1 Latent Feature Construction
We extract features directly from the latent domain of Stable Diffusion 2.1 (SD2.1). Each image $x\in\mathbb{R}^{3\times H\times W}$ is resized to $256\times 256$ and encoded through the pretrained SD2.1’s Variational Autoencoder (VAE) into a latent $z_{0}$ scaled by the constant $s$ as in
@@
where $\mathcal{E}(\cdot)$ is the VAE encoder. To reveal and aggregate residual diffusion traces, we reapply the forward diffusion process at discrete steps $t\in\mathcal{T}=\{0,5,10\}$, sampling
@@
where $t$ denotes the diffusion timestep and $(\alpha_{t},\sigma_{t})$ follow the cosine schedule with
@@
Each $z_{t}$ is normalized by $\sigma_{t}$ to maintain consistent scale, producing $\{z_{t}/\sigma_{t}\}_{t\in\mathcal{T}}$. We adopt the early-step set $\{0,5,10\}$ because these timesteps lie in the low-noise regime where generator-specific residuals remain most informative. This choice allows us both to study how the signal leak evolves across noise levels and to temporally aggregate its most discriminative components before they are suppressed by high-variance noise at larger $t$.
Figure 2: Interpretability of Proto-LeakNet (t-SNE Evolution). Layer-wise evolution of the embedding space, illustrating an approximate progression of how prototypes and attention progressively refine and generalize class separation across layers.
@@ 3.2 Discriminative Backbone
The feature tensors $\{z_{t}\}_{t\in\mathcal{T}}$ are encoded by a ResNet18 [he2015deepresidual] backbone $\phi(\cdot;\theta)$, with the first convolution adapted to match the channel dimensionality of $z_{t}$, where $\theta$ denotes the trainable backbone parameters. For each timestep $t$, the backbone produces a latent embedding
$h_{t}=\phi(z_{t};\,\theta)\in\mathbb{R}^{D},$ (4)
where $D$ is the embedding dimensionality. A learnable attention module assigns relevance to each timestep via
@@ 3.3 Prototype-Based Attribution Head
Each class $c\!\in\!\{1,\dots,C\}$ is represented by $M$ learnable prototypes, $p_{c,m}\!\in\!\mathbb{R}^{D}$, where $m\in M$, which serve as representative points in latent space. Empirically, four prototypes yielded compact yet well-separated latent clusters. An attention feature-wise gating vector $w\!\in\!(0,1)^{D}$ is computed using a small MLP:
@@
where $\tau>0$ is a learnable scalar that determines the aggregation smoothness. Posterior probabilities are computed as
@@
where $B$ is the mini-batch size and $y^{(b)}$ are ground-truth labels. All parameters $\{\theta,A,b,p_{c,m}\}$ are optimized jointly with AdamW and weight decay. Supplementary Material will further detail such parameters.
@@
The most activated prototype $\arg\max_{m}\pi_{c,m}$ identifies the latent region that best matches the input. Together, the feature gates $w$, temporal weights $\{a_{t}\}$, and prototype responsibilities $\{\pi_{c,m}\}$ provide a three-level interpretability hierarchy visible in Fig. 2 in detail with the effective distribution.
@@ 3.4 Representation-Level Generalization via Density Estimation
$p_{\text{KDE}}(h)=\frac{1}{N_{\mathrm{c}}(2\pi\sigma^{2})^{\frac{D}{2}}}\sum_{i=1}^{N_{\mathrm{c}}}\exp\!\left(-\frac{\|h-\bar{h}^{(\mathrm{c})}_{i}\|^{2}}{2\sigma^{2}}\right)$
@@
4 Dataset and Evaluation Metrics
Our evaluation builds on WILD [bongini2025wild], a 20k-image benchmark comprising high-quality samples from a closed set of text-to-image generators and an open set of additional state-of-the-art models, with no prompt or model overlap. WILD is selected for its realistic, artifact-free images, preventing attribution bias rooted in visible flaws. The dataset also includes post-processed variants introducing realistic degradations.
Closed set (10k images). The closed set includes ten text-to-image models: Adobe Firefly [firefly], DALL·E 3 [dalle3], FLUX.1 [flux2024], FLUX 1.1 Pro [fluxpro], Freepik [freepik], Leonardo AI [leonardoai], Midjourney [midjourney], Stable Diffusion 3.5 Large [stable_diffusion_35_large], Stable Diffusion XL Turbo [sdxl_turbo], and StarryAI [starryai]. A pool of 1,000 prompts is applied uniformly across all generators. The official split (also used for our experiments) contains 5,000 training images, 2,000 validation images, and 3,000 test images, with splits defined at the prompt level.
Post-Processed Closed-Set. The closed-set post-processed test images underwent 1 (Step 1), 2 (Step 2), and 3 (Step 3) random transformations selected from compression, cropping, resizing, rotation, blur, photometric changes, grayscale conversion, and super-resolution.
Open set (10k images, test-only). The open set includes ten additional generators, again with 1,000 images each: DALL-E, 1 [ramesh2021zeroshot], DeepAI [deepai], HotpotAI [hotpotai], NVIDIA Sana [xie2024sana], Stable Cascade [pernias2023wuerstchen], Stable Diffusion Attend&Excite [sd_ae], StyleGAN [karras2019stylebasedgeneratorarchitecturegenerative], StyleGAN2 [karras2020analyzingimprovingimagequality], StyleGAN3 [karras2021alias], and Tencent Hunyuan [li2024hunyuandit]. This set includes GAN models and text-to-image diffusion/transformer models not found in the closed set. All open-set images were used exclusively for testing.